OpenStudy (anonymous):

x2-6x-7 polynomial expressions ?

OpenStudy (gudden):

@youngstar10 .. do u want to factorize it for the zeros..?

OpenStudy (anonymous):

I have to use the quadratic formula

OpenStudy (gudden):

so do u know the quadratic formula..? if yes, then try to solve it..

OpenStudy (anonymous):

I know its x=-b+or - the square root of b^2-4ac divided by 2a

OpenStudy (gudden):

yeah it is..

OpenStudy (gudden):

The general formula of a quadratic equation is \[ax^{2} + bx + \]

OpenStudy (gudden):

c

OpenStudy (gudden):

so try using these..in the equation ... or should i solve it , first and then explain..?

OpenStudy (anonymous):

solve it then explain im a visual learner

OpenStudy (gudden):

\[d = \sqrt{b ^{2}- 4ac}\] => D = sqrt 36 + 28 = \[\sqrt{64} = \pm 8\]

OpenStudy (gudden):

and.. u know the value of b = -6 => -b = 6 and a = 1.. so using it in the equation we get first zero to be = (6 +8)/2 = 7 and the other zero to be = (6-8)/2 = -1